### v1

# SKYSURF IX - The Cosmic Optical and Infrared Background from Integrated Galaxy Light Measurements

Abstract
As part of the SKYSURF Hubble Space Telescope (HST) Legacy Archival program we present galaxy number counts which yield measurements of the extragalactic background light (EBL) at 15 different wavelengths. We have processed 82,752 HST images across 23 filters into 16,686 mosaics using the same software and processing pipeline throughout. Using 17/23 filters that give reliable galaxy counts, we constrain the integrated galaxy light (IGL) with a 1.5-9% error between 0.3 and 1.6 m in combination with 8 bands from WAVES (Wide Area VISTA Extragalactic Survey) and DEVILS (Deep Extragalactic Visible Legacy Survey). While HST was never intended to undertake large area surveys, through extensive quality control and filtering, we were able to extract a reliable and representative sample of fields distributed across the sky. Our final catalogs cover a combined , with individual filters covering areas ranging from . The combination of numerous independent sight-lines and area coverage allows us to reduce cosmic variance uncertainties in deep number counts to 0.06%-1.8%. For the first time we are able to establish a measurement of the IGL, , at 0.59 m using HST data. We obtain a cosmic optical background value of . Different techniques used to measure the COB, both directly and indirectly, have recently converged indicating that the COB arises almost exclusively from processes within galaxies. This in combination with the recent values reported from New Horizons and very high energy (VHE) constraints leaves very little room for any diffuse emission coming from outside the Milky Way.
1 Introduction
The extragalactic background light (EBL) is composed of all radiation originating from outside of our galaxy (Hill et al., 2018; Driver, 2021). The EBL is typically broken down into wavelength regimes such as the cosmic optical background (COB ) , cosmic infrared background (CIB ) and cosmic radio background (CRB); e.g. (Hill et al., 2018). The cosmic microwave background (CMB) is distinct from other astrophysical backgrounds as it is a relic from when the early Universe cooled to a temperature where neutral atoms could form. The remaining EBL from -ray to radio wavelengths contains information about all physical processes that have produced photons since recombination and thus can be considered separate from the CMB. The COB and CIB dominate the EBL and provide roughly equal energy contributions (Driver et al., 2016b; Koushan et al., 2021), though they both respectively fall a factor of 40x below the intensity of the CMB. The COB is perhaps the most important component of the EBL due to its intimate relationship with the cosmic star formation history (CSFH). The CSFH is defined as the volume averaged rate at which stars form as a function of either look-back time or redshift, for a through review see (Madau & Dickinson, 2014). The COB can be directly predicted from the CSFH assuming an implementation of a stellar initial mass function (IMF), dust attenuation model, metallicity evolution history, and stellar population model which has been demonstrated in works such as Andrews et al. (2017) and Koushan et al. (2021). Thus, empirical constraints of the COB can be used to test these models. The contribution from AGN is also important; especially at UV and mid-IR wavelengths, and is also able to be modeled. The cosmic AGN luminosity history (CAGNH) is known to co-evolve with the CSFH (D’Silva et al., 2023). Despite its importance, the COB remains one of the most difficult portions of the EBL to study in large part due to the presence of foregrounds which outshine the EBL, see (Bernstein et al., 2002; Windhorst et al., 2022; Caddy et al., 2022; Carleton et al., 2022; Windhorst et al., 2023).
Chief among these foregrounds is the Zodiacal light (ZL) that originates from sunlight scattered off of diffuse dust grains within the inner solar system. The ZL is the brightest of the optical and NIR foregrounds and constitutes as much as of the sky brightness at (Windhorst et al., 2022). It is also time variable, which requires observations of diffuse light over long periods of time from outside of the Earth’s atmosphere as done in Kelsall et al. (1998). More recent studies of the ZL such as Korngut et al. (2022) show that the theoretical model posed in Kelsall et al. (1998) still provides an accurate description of the ZL when allowing for a free offset parameter. Part of the SKYSURF project was designed to improve the models of the ZL and to attempt to place upper limits on the EBL at NIR wavelengths using HST, see, in particular, O’Brien et al. (2023). However, from low Earth orbit (LEO), stray-light from the Sun and reflections from the Earth and moon pose significant challenges to surface brightness studies using HST. In Caddy et al. (2022), the authors show that at optical wavelengths scattered light can be even more intense than the ZL when HST is pointing close to the Earth’s limb, as it approaches the daylight portion of its orbit. The final diffuse foreground is known as the diffuse galactic light (DGL). The DGL is produced from starlight scattered by the interstellar medium of the Milky Way and cannot be avoided when attempting a direct measurement of the EBL. Subtracting these foreground sources of diffuse light is difficult but necessary to obtain a direct measurement of the EBL (Bernstein et al., 2002; Bernstein, 2007).
Efforts to disentangle the DGL from other foregrounds include the dark cloud method employed by works such as Mattila et al. (2017), in which an optically thick molecular cloud is used to establish lines of sight which end at the cloud or continue into extragalactic space. However, results may be impacted by scattered starlight from the cloud. Recent attempts to measure the EBL directly from the outer Solar System have also been performed, e.g., Lauer et al. (2022) and Postman et al. (2024) in which the New Horizons mission was able to take optical images data using the LORRI camera at distances greater than 50 AU from the Sun, which should place the spacecraft well beyond the interference produced by the Zodiacal light, Earth, and Moon. Efforts were made to avoid bright stars and correct for both DGL and scattered starlight off of the spacecraft.
Indirect detection techniques have become possible with advancements in the ability to study high energy -ray sources. The interaction between photons from distant -ray sources, typically blazars, and optical EBL photons results in an attenuation of TeV energy -rays through particle/anti-particle pair production interactions as discussed in Desai et al. (2017). Using blazar spectra with established redshifts, they are able to estimate the attenuation of -ray spectra by EBL photons along the line of sight between the source and detectors. While difficult, this method (hereafter VHE) is less susceptible to foregrounds and is also unaffected by issues such as cosmic variance and zero point uncertainties, which plague the IGL. However VHE techniques typically apply a universal spectral shape to the -ray sources and infer deviations from this spectrum as entirely due to interactions with EBL photons or intergalactic magnetic fields. The technique typically requires an existing COB model and just constraints the normalization of thhe model. Recent developments in VHE studies of the EBL, (e.g. Gréaux & Biteau, 2023; Gréaux et al., 2024) have adopted a Bayesian framework to avoid limitations introduced by assuming a model and spectral shape for both the EBL and gamma ray source. Using the observed -ray spectra of blazars they present preliminary results in which they now extract the COB shape and amplitude. Uncertainties in measurements of the COB established by VHE studies to 15% and will be developed further in the coming years.
A commonly employed method to place constraints on the EBL without accounting for foregrounds is to use galaxy number counts and measure the total intensity of all resolved galaxies in works such as (e.g. Madau & Pozzetti, 2000; Totani et al., 2001; Xu et al., 2005; Dole et al., 2006; Hopwood et al., 2010; Keenan et al., 2010; Berta et al., 2011; Béthermin et al., 2012; Driver et al., 2016b; Koushan et al., 2021; Windhorst et al., 2022, 2023, Carter et al. 2025(submitted)). The IGL can be well constrained by surveying a large enough volume to obtain a representative sample of the Universe. Galaxy number counts, or rather the galaxy flux density, can be integrated to provide a measurement of their combined intensity. While establishing robust lower limits, the IGL is prone to uncertainties introduced by cosmic variance and magnitude zero-point differences present between the facilities providing the number counts. Cosmic variance and zero point errors remain the dominant sources of uncertainty in measuring the IGL. In work such as Driver et al. (2016b) these contribute uncertainties of 10% and 5%, respectively. More recent studies such as Koushan et al. (2021), through the use of larger area surveys and updated zero point calibrations, were able to bring these uncertainties down to 4%. However, Driver et al. (2016b) and Koushan et al. (2021) relied only on a small number of deep HST fields. Work such as Windhorst et al. (2011) and Rafelski et al. (2015) provide the faintest optical galaxy number counts and have only been surpassed by JWST recently in the NIR (Windhorst et al., 2023). The two HST surveys included in Driver et al. (2016b) and Koushan et al. (2021) cover a combined 40-60 arcmin2, and thus cosmic variance remains the dominant source of uncertainty in these surveys at 10-15% as estimated using the formula developed in Driver & Robotham (2010). Uncertainties in the IGL have also been reduced to between by Carter et al. 2025 (submitted). SKYSURF aims to use a significant fraction of the area covered by HST over 20 years of observations to further reduce uncertainties in measurements of the IGL primarily by reducing the cosmic variance errors of individual HST deep surveys used in previous works.
Despite efforts to measure the COB both directly and through the IGL, disagreements between the upper limits established by work of Carleton et al. (2022) O’Brien et al. (2023), Mattila et al. (2017), Lauer et al. (2022) and lower limits established by the IGL in Driver et al. (2016b) and Koushan et al. (2021) remained factors of apart. The newest direct measurements of Postman et al. (2024) from New Horizons suggest that the tension between the IGL and direct measurements has finally been resolved. Again using the LORRI instrument aboard New Horizons, measurements have yielded a significant detection of the EBL at the LORRI pivot wavelength of , with the main change in Postman et al. (2024) being an improved method for estimating the DGL. Within their respective errors, the measurements of Postman et al. (2024), Koushan et al. (2021), and preliminary VHE measurements from Gréaux & Biteau (2023) have finally converged. The explanation posed in Postman et al. (2024) is that the IGL composes the most significant component of the COB and the majority of extragalactic optical flux originates from within galaxies. The JWST work of Windhorst et al. (2023) suggests that no more than 20% of the COB can originate from outside of galaxies. As such, efforts to reduce uncertainties in the IGL can better constrain models and provide a near complete census of photon production from baryon-photon decoupling until the present day.
In this paper, we begin by introducing the WAVES (Wide Area Vista Extragalactic Survey) photometric data, though for a complete discussion of the WAVES survey see the WAVES overview paper(Bellstedt et al. (in prep)). For our other supplementary data set we provide a brief overview of the DEVILS (Deep Extragalactic VIsible Legacy Survey) photometric data (Davies et al., 2021). We then detail the processing of the SKYSURF database from assembly, reduction, and data quality checks. The production of galaxy number counts and the necessary steps to reduce HST object catalogs to usable number counts is discussed in Section 3. In Section 4 we discuss the most significant issue posed by using HST to produce wide area counts given HST’s bias towards observing cluster fields. In Section 5 we present measurements of the IGL, sources of uncertainty, and how they can be improved. We then discuss the difficulties imposed by remaining bias in our HST number counts. Finally, we compare our measurements to existing measurements of the COB from IGL studies, VHE constraints, and direct measurements in Section 6. Throughout we assume a standard 7-3-7 cosmology with H0 = 70 , and .
2 Data and Data Reduction
The data presented in this work is derived from catalogs and/or galaxy number counts as part of the WAVES (Bellstedt et al. in prep), DEVILS (Davies et al., 2021), and SKYSURF (Windhorst et al., 2022) survey programs. Here we introduce the relevant surveys and present our SKYSURF HST data processing pipeline.
2.1 The WAVES Survey
WAVES is a spectroscopic redshift survey to be conducted using the European Southern Observatory’s 4-Metre-Multi-Object Spectroscopic Telescope (4MOST) (de Jong, 2011). In this work, bright galaxy number counts are produced in preparation for the WAVES survey (Driver et al., 2016a, 2019, Bellstedt et al. in prep) which covers a region of the sky. The galaxy catalogs used for WAVES target selection (Bellstedt et al. in prep) are produced from imaging taken from the VISTA KIlo-degree Infrared Galaxy (Edge et al., 2013, VIKING) and the VST Kilo-Degree Survey (Kuijken et al., 2019, VST-KiDS), both of which are ESO public data sets. The data was WCS aligned using ProPane (Robotham et al., 2023), and object detection is done using an rizY inverse variance stack using ProFound (Robotham et al., 2018). All bright galaxies were inspected for fragmentation following the detection phase and fixed. Following the detection phase, fluxes in each band were determined using forced aperture photometry from the detection stack. Stars brighter than GAIA g mag were masked. Star-galaxy separation was achieved using an unsupervised machine learning technique developed in Cook et al. (2024). See Bellstedt et al. (in prep) for full details on these steps. In this work we use WAVES input catalog v1.2 with 7 bands (ugrizYJH) that overlap with the HST filter system. The completeness limit of each band is shown in Table 1 where completeness limits were defined by stepping back by one half-magnitude bin from where the counts reach a maximum.
2.2 The DEVILS Survey
For intermediate magnitude number counts we turn to DEVILS (Davies et al., 2021), using all three DEVILS fields, D02, D03, and D10. The three fields are centered on well-studied regions, XMM-LSS, ECDFS, and COSMOS, respectively. The DEVILS galaxy number counts are complete to AB 24 mag across all 3 fields and cover a combined area of 6 . A brief summary of areas and completeness limits are provided in Table 1. We elect to use the same 7 bands chosen from the WAVES survey, i.e. , . We note that D02 and D10 were observed by the Hyper Suprime Cam (HSC) (Aihara et al., 2017) in griz with the YJH bands observed by VISTA. In D03 the griz imaging was obtained by the VST. Finally, u-band imaging was obtained from the Canada-France-Hawaii-Telescope (CFHT) and VST.
Source detection was done using ProFound on an inverse variance weighted stack of three photometric bands, YJH. This was done to maximize depth and to avoid any potential issues caused by artifacts in a single band. Source detection was performed on this stack and the resulting object contours were used for flux extraction in all bands. Star-galaxy separation was done using the relationship between half-light radius and Y band magnitude, as well as a relationship between magnitudes from the Y, J, H, and K bands. For a complete discussion of the DEVILS survey, see Davies et al. (2021) and DEVILS data-release (Davies et al. 2025, Submitted).
Band | WAVES Area | WAVES Completeness Limits | DEVILS Area | DEVILS Completeness Limits |
ABmag | ABmag | |||
u | 1067.86 | 9.75 - 23.25 | 5.96 | 15 - 25.5 |
g | 1067.86 | 9.75 - 23.25 | 5.96 | 15 - 25 |
r | 1067.86 | 10.75 - 23.25 | 5.96 | 15 - 25 |
i | 1067.86 | 9.75 - 22.25 | 5.96 | 15 - 24.5 |
z | 1067.86 | 10.25 - 22.25 | 4.48 | 14.5 - 24.5 |
Y | 1067.86 | 9.75 - 21.75 | 5.97 | 15 - 24.5 |
J | 1067.86 | 9.75 - 21.25 | 5.96 | 15 - 24 |
H | 1067.86 | 9.24 - 21.25 | 5.96 | 14.5 - 24 |
2.3 The SKYSURF HST Archival Program
SKYSURF is an HST cycle 27-29 Legacy Archival program (AR-09955 & AR-15810) designed to use a significant amount of HST imaging data to measure the sky-surface brightness (sky-SB) across the wavelength range covered by HST, i.e., . This requires the reprocessing of archived HST data using consistent software pipelines throughout. During the last two years, the SKYSURF project has measured the sky-SB and divided it into its components, the ZL, DGL, and EBL, (e.g. O’Brien et al., 2023; Carleton et al., 2022; Windhorst et al., 2022, Carter et al. 2025 (Submitted)). The SKYSURF project is split into two separate teams using independent processing pipelines and methodologies to obtain mosaics, sky-SB values, and object catalogs. This paper constitutes the portion of the project led from the University of Western Australia focused on integrated galaxy light (IGL) measurements using object catalogs generated using the ProPane and ProFound software packages, (Robotham et al., 2023, 2018) as was used by WAVES and DEVILS. In Carter et al. 2025 (submitted), an independent processing of the SKYSURF archive was performed using the AstroDrizzle (Gonzaga et al., 2012) for image alignment and stacking. Source detection was done using SourceExtractor (Bertin & Arnouts, 1996). Our overriding objective is to produce reliable galaxy number counts across as broad a magnitude and wavelength range as possible. We begin our SKYSURF selection by only including HST imaging with an exposure time of 200 seconds or longer and at galactic latitude . A projection showing the location of SKYSURF data used in this work is shown in Figure 1. All data was downloaded from the Mikulski Archive for Space Telescopes (MAST)111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html.
Filter | # of Fields | Final Area () |
ACSWFC F435W | 186 | 0.564 |
ACSWFC F475W | 335 | 0.992 |
ACSWFC F555W | 58 | 0.175 |
ACSWFC F606W | 978 | 2.940 |
ACSWFC F625W | 188 | 0.587 |
ACSWFC F775W | 424 | 1.350 |
ACSWFC F814W | 2201 | 7.04 |
ACSWFC F850LP | 299 | 1.04 |
WFC3UVIS F336W | 97 | 0.190 |
WFC3UVIS F390W | 80 | 0.160 |
WFC3UVIS F606W | 279 | 0.529 |
WFC3UVIS F814W | 479 | 0.880 |
WFC3IR F105W | 345 | 0.460 |
WFC3IR F110W | 245 | 0.329 |
WFC3IR F125W | 344 | 0.531 |
WFC3IR F140W | 288 | 0.466 |
WFC3IR F160W | 1007 | 1.372 |
2.4 SKYSURF Data Reduction
We use the data assembled in SKYSURF-1 Windhorst et al. (2022) initially filtered as described in Section 2.3 and construct the mosaics summarized in Table 2 as follows. Starting with calibration level 3 products from MAST we produce sky-subtracted science images, inverse variance maps, and masks for each image extension using the ProFound image processing software (Robotham et al., 2018). A flow chart summarizing the data processing steps is shown in Figure 2. In this work, we use the term ”mosaic” to refer to any processed data regardless of how many exposures or observing programs are included. For example, some SKYSURF final data products consist of a single exposure while others are built from hundreds of exposures. SKYSURF mosaics comprised of multiple overlapping observing programs are referred to as multi-visit products. Mosaics produced from single observing programs are referred to as single-visit products. A series of footprints showing both complex and simple mosaic tiling patterns are displayed in Figure 3.
To ensure accurate alignment during stacking of multi visit data we further restrict our input images to those recalibrated to GAIA DR3 WCS. For mosaics composed only of a single observing program, we choose the data with the original WCS system associated with the program at the time of observation in order to be maximally inclusive. An example intermediate data product is shown in Figure 4 with a final mosaic shown in Figure 5. Mask generation is done using the HST data quality (DQ) FITS extension. We mask pixels associated with potential cosmic rays, and dilate these regions by three pixels to be conservative in removing potentially contaminated pixels not flagged in the original DQ extension. All other bad detector pixels and readout columns are masked but not dilated. Each individual sky subtracted image, inverse variance map, and mask are used to produce inverse variance weighted mosaics using the propaneStackWarpInvar function in ProPane (see Figure 4). The science mosaics are in flux units of electrons per second. Each mosaic contains an image extension, inverse variance map, weight map, and exposure time map (see Figure 5). The processing software pipeline used for the project is open source and publicly available on github222github.com/stompkins7192/HST_Builder.
3 Object Catalog and Counts Construction
Here we discuss the methods used for the SKYSURF catalog construction. Initial object detection was achieved using ProFound (Robotham et al., 2018). We first used a subset of images from the database with a range of filters and exposure times to optimize the settings in ProFound for object detection via an iterative process of varying settings and inspecting the resulting diagnostics, checking for completeness, reliability, and optimal deblending. The final profoundProFound object detection settings are shown below.
profoundProFound ( image = sci,
header = sci[,]header,
mask = mask,
plot = FALSE,
magzero = magzero,
pixscale = ps,
rem_mask = TRUE,
keepim = FALSE,
skyRMS = (invar)-0.5,
#Use inverse variance map as skyRMS
skycut = 1.5,
pixcut = 4.0,
ext = 4,
smooth = TRUE,
sigma = 0.8,
tolerance = 8,
reltol = 1,
cliptol = 300,
size = 5,
iters = 7,
box = c(((dim(mask)[1])/3),
((dim(mask)[2])/3)),
boxiters = 2,
roughpedestal = FALSE
pixelcov = FALSE
boundstats = TRUE,
nearstats = TRUE,
redosky = TRUE,
redoskysize = 21,
SBdilate = 2 )
We then applied profoundProFound, the top level source detection and cataloging function in ProFound, to all 16,686 mosaics including both single and multi-visit products and detected a total of 153 million objects (including stars, galaxies, and artifacts). All magnitude values for galaxy number counts used in this work are sky-subtracted apparent AB magnitudes as calculated by profoundProFound. After running source detection, we implement a four step process to go from our raw catalogs to galaxy number counts. These steps are galactic extinction corrections, star-galaxy separation, bright star masking, and completeness corrections as described in the following sections.
3.1 Galactic Extinction Corrections
We correct for galactic extinction using the Planck satellite dust maps accessible through the web tool333https://irsa.ipac.caltech.edu/applications/DUST/ with extinction estimates provided by Finkbeiner et al. (1998). Given the small angular size of SKYSURF fields compared to the resolution of the dust maps, we elect to use the center of the field to obtain a single E(B-V) value and thus apply one extinction value to all objects in that field. Using the Spanish Virtual Observatory tool provided by (Rodrigo & Solano, 2020)444http://svo2.cab.inta-csic.es/theory/fps/, we obtain values for each HST filter. The Av values adopted for each HST filter used in this work along with the pivot wavelengths adopted in this work are listed in Table 3. In Figure 6 we show the Galactic extinction values for each mosaic as a function of Galactic latitude. By design, SKYSURF omitted all observations within 10 degrees of the galactic plane, as indicated by the vertical lines in Figure 6.
Filter | Pivot Wavelength (Å) | Value |
ACSWFC F435W | 4369.62 | 1.32 |
ACSWFC F475W | 4783.06 | 1.20 |
ACSWFC F555W | 5413.32 | 1.04 |
ACSWFC F606W | 5962.2 | 0.924 |
ACSWFC F625W | 6323.81 | 0.853 |
ACSWFC F775W | 7763.11 | 0.645 |
ACSWFC F814W | 8102.91 | 0.602 |
ACSWFC F850LP | 8945.58 | 0.505 |
WFC3UVIS F336W | 3371.04 | 1.69 |
WFC3UVIS F390W | 3955.42 | 1.47 |
WFC3UVIS F606W | 5946.48 | 0.93 |
WFC3UVIS F814W | 8107.52 | 0.603 |
WFC3IR F105W | 10542.47 | 0.395 |
WFC3IR F110W | 11696.34 | 0.34 |
WFC3IR F125W | 12503.91 | 0.30 |
WFC3IR F140W | 13983.36 | 0.256 |
WFC3IR F160W | 15437.71 | 0.22 |
3.2 Star-Galaxy Separation
Without color information for all SKYSURF fields, only object sizes and single filter magnitudes can be used for star-galaxy classification. For star-galaxy separation we use the apparent magnitude and R50 parameters provided by ProFound. R50 is defined as the radius in arcseconds within which half of an object’s integrated flux is contained. Figure 7 shows the apparent magnitude versus plane for the ACSWFC F606W filter. The stars are shown as blue points and follow a nearly constant versus magnitude relationship below AB 20 mag. Between AB 20-22 mag, we fit a double Gaussian model in to the histogram of as shown in Figure 8 with the dividing boundaries marked. The stars form a narrow locus until AB 20 mag when their PSF begins to over-saturate the detector and their measured sizes increase. We use the width of the normal distribution occupied by the stars to define a gentle slope starting from AB 20.25-21.75 mag and ending at the mag boundary at which the populations blend together at faint magnitudes. While the minimum size between the two populations shown in Figure 8 occurs at -0.8, visual inspections of objects fainter than AB 21.25 mag confirm that objects between the displayed boundary and the galaxy locus confirmed that these objects are galaxies. Star-galaxy separation is done before performing galactic extinction corrections. At bright magnitudes we define a steeper slope from the size-magnitude relationship of the stars that forks upwards when the stars begin to over-saturate the detector. This is done independently for each filter. We also use the measured width of the stellar population, , to define a boundary between real objects and artifacts, which are denoted by green points in Figure 7.
Visual inspections suggest that objects with sizes less than below the median value of the stellar locus are artifacts and thus we exclude them for all purposes in this work. The truncation point of star galaxy separation is derived by attempting to fit a double-Gaussian model to the parameter in each half-magnitude bin. Once the double Gaussian model is no longer successful due to the blending of the populations we stop attempting to separate stars from galaxies. This typically occurs at AB mag for ACSWFC and WFC3UVIS. This occurs at slightly brighter magnitudes, mag, for WFC3IR given its lower resolution. Note that at this flux level the galaxy number density far outnumbers that of the stars. We also choose to mask all objects brighter than AB 15 mag given the risk of fragmentation during source detection. The results of these selections are shown in Figure 7 with our galaxy sample shown in red, stars and objects brighter than AB 15 mag shown in blue, and artifacts shown in green.
3.3 Bright Star Masking
Manually inspecting fields and their associated segmentation maps also revealed that ProFound identifies diffraction spikes and over-saturated regions associated with bright stars as objects. To combat this problem and flag potential diffraction spikes, we define a circular mask around stars within which all objects are flagged. The radius of this mask was determined by inspecting the ProFound segmentation maps and measuring the sizes of diffraction spikes detected by ProFound for stars of different magnitudes.
The results are shown in Figure 9, where our derived relationship is , where M is the size mask radius in pixels as a function of apparent magnitude with slope, a, and y-intercept, b.
This formula is used to define a circular region around stars to mask and exclude. Objects falling within these regions are flagged in the catalog as they could be diffraction spikes or have their photometery affected by the presence of the nearby star. To adjust for this in post processing, we re-calculated the area of the image after drawing circular masks around flagged objects and removing the enclosed pixels from the area calculation for each frame. We use the apparent magnitudes as measured by ProFound in this process (). Stars with mag are faint enough that ProFound does not deblend diffraction spikes into secondary objects and the star does not significantly affect the photometry of adjacent objects. We performed this for the WFC3IR F160W, WFC3UVIS F814W, and ACSWFC F814W filters, with the values for slope, a, and y intercept, b, shown in Table 4. The resulting relationships between mask size and magnitude are used for all other filters in the corresponding instrument (WFC3IR, WFC3UVIS, and ACSWFC). Using a well-studied and defined HST point-spread-function mask is not feasible for SKYSURF as in regions where frames taken at different orientations are stacked, multiple sets of diffraction spikes and other associated issues are seen as in Figure 10.
Filter | Slope (a) | Y-intercept (b) |
ACSWFC F814W | -0.165 | 4.74 |
WFC3UVIS F814W | -0.146 | 4.59 |
WFC3IR F160W | -0.137 | 3.99 |
3.4 Completeness Corrections
The nature of SKYSURF means that different regions in the same mosaic reach different magnitude limits as exposure times can vary significantly within individual mosaics. Thus, completeness limits for object counts cannot be defined by selecting the peak of the combined number counts within a mosaic. Without accounting for exposure time, the counts would gradually roll over (flatten) as sub-regions within their mosaic reach their completeness limits. We therefore developed a strategy using exposure time maps for each mosaic as shown in Figure 11. We select exposure time bins which increase in width by a factor of 2 beginning with 200 seconds (the minimum exposure time for frames comprising the SKYSURF database).
For each exposure time bin, we calculate the combined area across all single & multi visit mosaics and generate combined object counts for each exposure time region. Only exposure time regions covering more than 10% of the unmasked area in a mosaic are considered for object counting purposes. We perform this operation in catalog space by adding the exposure time for each object as an additional column along with the unmasked area associated with that specific sub-region. We then collate objects with similar exposure times in the catalog and sum the pixel area from the final catalog containing all fields. Figure 11 shows an example mosaic demonstrating this process from the WFC3UVIS F814W filter. In Table 5 we show the corresponding completeness limits, area coverage, and completeness limits for the entire WFC3UVIS F814W filter mosaic set.
Exptime Range (s) | Median Exptime (s) | Area | Fraction | Completeness Limit (ABmag) |
400-800 | 780 | 0.154 | 0.159 | 24 |
801-1600 | 1313 | 0.172 | 0.177 | 24 |
1601-3200 | 2508 | 0.460 | 0.475 | 24.5 |
3201-6400 | 4931 | 0.135 | 0.139 | 25 |
6401-12800 | 7420 | 0.022 | 0.023 | 26 |
12801-25600 | 18690 | 0.021 | 0.0214 | 26 |
25601-51200 | 27250 | 0.006 | 0.006 | 27 |
4 Addressing HST’s Observing Biases
While the SKYSURF database covers a significant area, this area is not necessarily representative of the Universe at large. Cluster and nearby galaxies compose a significant fraction of the observations performed by HST and thus SKYSURF mosaics are prone to significant bias. We implement three strategies to address this. First we flag and ignore observing programs targeting known globular clusters or nearby galaxies by using the target name from the FITS header. Targets with names such as M31, 47TUC, and NGC were rejected. However, many proposals do not provide a target name and thus this is not the only method employed for sorting data.
Secondly, we turn to statistical methods to remove dense cluster fields and under dense fields with either processing errors or due to observations of targets such as nebulae or nearby large galaxies that occupy a significant portion of the field of view. To remove such fields, we first create galaxy number counts from the catalog for each filter. We then perform a zero point correction to move both WAVES and DEVILS data to the chosen HST filter. This is done using the FilterTranBands function in ProSpect software package (Robotham et al., 2020). For a detailed discussion of the magnitude zero-point corrections see Section 5.3.
From this we fit a spline to the number counts and the energy curve, see the dashed line fit in Figure 12 (centre). This represents our reference best fit line for that filter. The energy curve shows the flux density contained in each galaxy number count bin. From AB 20-24 mag we take the difference between the reference fit to the energy curve and the energy curve data points for each individual SKYSURF field. The mean value of this quantity for each field is recorded and shown in the histograms in Figure 12 (bottom). A 3 clip is performed to remove outlier fields, and the entire process is repeated until the sigma clipping converges when no additional fields are clipped. Figure 12 (centre) shows the effectiveness of the sigma clipping. With the left panel showing a disagreement with the WAVES data but agreement after removing over-dense fields. We also discard SKYSURF number counts at magnitudes brighter than AB 19.75 mag as brighter counts are far more robust from WAVES. However, for the F140W filter the entire magnitude range is used given the significant magnitude zero-point corrections between these filters and the WAVES ground-based Y&J band data.
Finally, HST data is also prone to over-densities caused by excessive fragmentation of nearby galaxies during source detection and especially nearby low surface brightness objects, dwarf galaxies, globular clusters, and in some instances faulty readouts. While most of the area in such fields remains usable, these fields may contain unrealistic object densities at faint magnitudes. We thus elect to perform a third filtering step. We measure the number of objects brighter than AB 26 mag classified as galaxies in subregions (400x400 pixels for ACSWFC and WFC3UVIS, 100x100 for WFC3IR. The highest object density subregion for each field is stored. We removed fields containing subregions more dense than that 99th percentile of the total subregion density distribution. An example of a problematic field which we are able to identify this way is shown in Figure 13.
5 Galaxy Number Counts and the IGL
Here we discuss and present the SKYSURF galaxy number counts for the HST filter sets used in this analysis. In Figure 14 we show the galaxy number counts (left), integrated galaxy light (centre), and the normalized sum of the IGL (right) for the F606W filter as an example. To convert from galaxy number counts to a measurement of the EBL, we use the following standard equations
The output of Equation 1 is then multiplied by the following constants to obtain the standard EBL units.
where is the number of galaxies in a magnitude bin between m and .
The best spline fit to the data in Figure 14 is shown as a dashed black line alongside a subset of 100 Monte-Carlo spline fits to the data as described in Section 5.1. The splines shown in Figure 14 use 10 degrees of freedom. We report an extrapolated IGL value where we integrate between AB 100 mag, which is indistinguishable from integrating between . These values are reported in Table 6. We also report lower limits in Table 7 where we integrate to the faintest complete magnitude bin. When our galaxy number counts do not reach faint magnitudes, and/or the slope of the faint number counts is poorly defined, there can be significant differences between the extrapolated and non-extrapolated IGL. For example, in the F625W filter where the extrapolated IGL value of , and the lower limit ie . We thus elect to report a conservative mean value in with an uncertainty stretching between the on the respective bounds. These values are given in Tables 8 & 11.
In this work we face four sources of uncertainty which affect the IGL. Poisson statistical error, cosmic variance (CV), magnitude zero point differences between bandpasses, and fitting errors. We now discuss each of these errors in detail below.
5.1 Poisson Statistical Errors
For all spline fitting discussed in this section we use a weight on each data point where is the typical (where N is the number of galaxies in each half magnitude bin) Poisson statistical uncertainty on each data point. Integration under the spline fits to the central panel in Figure 14 allows us to obtain our IGL measurements. When deriving the Poisson error, the magnitude zero-point corrections provided in Tables 12 and 13 are used without any uncertainty in order to isolate the Poisson error. We independently shift each data point by its 1 uncertainty, then an error-weighted spline is fit to the perturbed data and the IGL value is obtained using the same integration process described above in Equations 1 & 2. This process is repeated 10,000 times where the median, 17th and 83rd percentiles of the resultant distribution are used to define a percentage uncertainty due to Poisson statistical error only.
5.2 Cosmic Variance Errors
Calculating a cosmic variance error for our HST number counts is non-trivial because the area changes with depth as seen in Table 5. As such, we calculate a fractional CV error for each SKYSURF galaxy count bin. This calculation is done using the cosvararea function in the celestial R software package555https://github.com/asgr/celestial/. The function takes a number of independent sight lines, area, and redshift range to calculate a volume and return the CV error as a percentage using the formulation from Driver & Robotham (2010). We assume each SKYSURF field is fully independent for the purposes of calculating the number of independent sight lines. We obtain the redshift range from the DEVILS redshift catalogs in the D03 and D10 fields. We take the 17th and 83rd percentile of the redshift distribution for objects from AB 18-24 mag as the inputs for cosvararea. This is done separately for each band in the ugrizYJH set given that the redshift distribution changes with wavelength. We obtain the CV uncertainty for both WAVES and DEVILS using the cosvararea function as well, where the value is constant regardless of magnitude given all magnitude bins cover the same area and number of independent sight lines.
We again perform Monte-Carlo analysis by shifting the number counts up/down according to their fractional cosmic variance errors again following a normal distribution with a mean of 1 and standard deviation of . As noted, the data from each survey (WAVES, DEVILS, and SKYSURF) are shifted systematically and independently for each data set. An error weighted spline is fit to the data and integrated. This process is repeated 10,000 times where the median, 17th and 83rd percentiles of the resultant distribution are used to define a percentage uncertainty due to CV error.
5.3 Magnitude Zero Point Corrections
Magnitude zero point errors will also affect the data systematically. We choose to color-correct number count data from WAVES and DEVILS onto the HST filter system which produces a systematic magnitude offset for both surveys. We calculate the color correction terms using the ProSpect filterTranBands function. As input, the function takes a filter, input filter, reference filter, and redshift range. The function returns the 3 terms used in the linear mappings. ProSpect comes with a long list of pre-calculated apparent magnitudes in each filter set which are used as inputs for the following equation
where , , and are terms returned by filterTranBands. is the apparent magnitude from the input filter, is the magnitude for a reference filter, and is the desired magnitude in the target filter. For example, when correcting WAVES Y-band data to HST’s WFC3IR F105W filter, the target, input, and reference filters are WFC3IR F105W, Y VISTA, and H VISTA, respectively. Using 1,000 of the ProSpect provided magnitudes, the final zero-point correction values and their uncertainties are provided in Tables 12 and 13.
Since we merge number counts from filters that exist in both ACSWFC and WFC3UVIS such as F606W, we take the mean of the two correction values and their respective uncertainties to correct ground based data to the HST filter system. We then take the correction and shift the WAVES and DEVILS number count magnitudes according to a normal distribution by uncertainties on the colour transformation parameters. We do this 10,000 times to obtain a spread of IGL values and use the 17th and 83rd percentiles to obtain an uncertainty, see Figure 15. The magnitude zero point corrections and their uncertainties are listed in the appendix.
5.4 Fitting Error
For all of the spline fitting in this work, we use a spline with 10 degrees of freedom by default. We choose to introduce a fitting error by fitting the data with splines of order 8-12 and taking the best fit IGL value for each order of spline. This is consistent with what has been done in previous works, see Driver et al. (2016b); Koushan et al. (2021). To stop the spline fits from diverging at the bright we ignore any galaxy number count bins with 50 or fewer galaxies. See Figure 22 (rows 3-4) where bright galaxy number count bins in WAVES-g band (green points) could introduce diverging spline fits. For the F140W (Figure 25, row 2) filter where no ground-based data is available we simply ignore the brightest two magnitude bins which otherwise cause the splines to diverge.
5.5 Further Discussion Of Uncertainties and Biases
We present our final IGL measurements below in Figure 16, where extrapolated IGL data is shown in red, complete with errors, with lower limits displayed as upward facing blue arrows. At most wavelengths where our faint number counts are well constrained, the lower limits on the IGL within the 1 uncertainties on the extrapolated IGL, suggesting that contribution to the optical IGL from galaxies fainter than AB 28.5 mag is small, (, see Windhorst et al. (2022). Where there is a significant difference, we attribute it to a poorly constrained slope in the faint number counts.
We attribute discontinuities in the IGL, such as the jump between the F555W and F606W at 0.54 & 0.59 m, to the choice of ground-based data sets used in conjunction with HST. The F555W filter falls between the g and r band data from DEVILS and WAVES and given its limited area coverage the IGL measurement for F555W is sensitive to the ground-based data chosen to accompany it. We elect to use g-band given it provides a lower uncertainty. Ideally, v-band data would be used alongside the F555W data though without access to wide area v-band number counts we are forced to use g-band data and accept a much larger uncertainty. Thus the significant increase in the IGL between 0.54 & 0.59 m can be attributed to the choice of reference data and should not be taken as an indication of any underlying physical process.
We test the assumption that our errors can be added in quadrature to avoid imprecision in our final error budget. We performed all of our error analysis techniques on the data simultaneously, and as expected the results are consistent with adding independent errors in quadrature. Across all bands, the mean contributions from zero-point, cosmic variance, Poisson, and fitting errors are 2.3, 2.0, 0.98, and 1.35 %, respectively. Though the most significant on average, zero-point correction uncertainties can be attributed to the choice to use 19 HST filters at 15 different wavelengths to sample the COB at more wavelengths.
As seen in Tables 6 & 7, HST filters such as the F475W and F625W at pivot wavelengths of 0.475 and 0.631, m, respectively, are designed to overlap with ground based g&r band facilities. As such the filter transform errors are very small. However, in the final SKYSURF database these filters cover significantly less area and thus their CV uncertainties are much larger. This is why we chose to zero-point correct ground based reference data sets to the HST filter system as this avoided us having to choose which 8 HST filters to use with the ugrizYJH filter set from WAVES & DEVILS. Cosmic variance is unavoidable, and one of the main goals of SKYSURF was to reduce CV errors in faint galaxy number counts by sampling a large area and number of independent sight lines when compared to previous surveys at similar depths. We note the small, but consistent CV uncertainty introduced by WAVES at bright magnitudes. Despite covering , the CV error from WAVES is still and is in part responsible for the cosmic variance error floor where it is used. For the F140W, at a pivot wavelengths of 1.39 m, only SKYSURF data is used and thus the cosmic variance uncertainty is lower than 1.5%. However, without the bright counts provided by WAVES, this introduces significant Poisson statistical error.
While the F140W SKYSURF IGL measurement offers a glimpse of the advantages provided by relying on a single facility to provide galaxy number counts, it has the largest uncertainty almost entirely due to Poisson statistical error. Without wide-area surveys to provide bright galaxy number counts, the small number of objects at bright magnitudes yield unavoidable statistical errors (see Figure 17). Similar to CV uncertainties, Poisson statistical errors can only be reduced by wider area surveys. Future data sets from missions such NASA’s Nancy Grace Roman space telescope will cover a broad range of magnitudes and can produce IGL measurements without the need to combine galaxy number counts from different facilities, thus eliminating magnitude zero-point correction uncertainties entirely.
HST Filter | Pivot Wavelength | IGL Value | Zero-Point | CV | Poisson | Fitting Error | Total Error | |
F336W | 0.335 | 4.05 | 2.47 | 3.10 | 1.18 | 1.47 | 4.39 | |
F390W | 0.392 | 4.94 | 0.96 | 4.52 | 3.22 | 1.10 | 5.25 | |
F435W | 0.432 | 5.85 | 1.94 | 3.46 | 0.72 | 1.32 | 4.45 | |
F475W | 0.475 | 5.77 | 0.21 | 0.91 | 0.95 | 2.63 | 2.84 | |
F555W | 0.536 | 6.30 | 1.99 | 8.31 | 1.01 | 2.75 | 9.03 | |
F606W | 0.591 | 9.07 | 2.75 | 2.51 | 0.95 | 0.51 | 3.87 | |
F625W | 0.631 | 9.14 | 0.21 | 2.97 | 2.05 | 0.05 | 3.29 | |
F775W | 0.769 | 11.48 | 0.84 | 3.32 | 1.22 | 2.29 | 4.29 | |
F814W | 0.810 | 11.28 | 0.63 | 1.33 | 1.36 | 4.34 | 4.79 | |
F850LP | 0.903 | 11.28 | 0.46 | 3.05 | 0.51 | 0.23 | 3.13 | |
F105W | 1.05 | 11.41 | 2.58 | 1.69 | 0.40 | 0.40 | 3.20 | |
F110W | 1.15 | 11.17 | 4.72 | 1.16 | 0.36 | 0.36 | 4.89 | |
F125W | 1.25 | 10.95 | 1.14 | 0.28 | 0.46 | 0.45 | 2.02 | |
F140W | 1.39 | 10.19 | N/A | 1.32 | 5.26 | 1.24 | 5.56 | |
F160W | 1.54 | 11.27 | 1.08 | 1.11 | 0.27 | 0.18 | 1.58 |
HST Filter | Pivot Wavelength | ABmag Range | IGL Value | Zero-Point | CV | Poisson | Fitting Error | Total Error |
ABmag | % | % | % | % | % | |||
F336W | 0.335 | 14.75-27.0 | 3.21 | 6.17 | 1.24 | 0.18 | 0.05 | 6.30 |
F390W | 0.392 | 14.75-27.0 | 3.20 | 1.40 | 1.27 | 0.17 | 0.03 | 1.90 |
F435W | 0.432 | 13.25-27.0 | 5.31 | 2.12 | 1.20 | 0.23 | 0.01 | 2.45 |
F475W | 0.475 | 13.25-26.5 | 4.83 | 0.24 | 1.07 | 0.20 | 0.03 | 1.11 |
F555W | 0.536 | 13.25-27.5 | 5.17 | 2.14 | 8.31 | 0.28 | 0.16 | 8.58 |
F606W | 0.591 | 12.25-28.5 | 8.09 | 2.79 | 1.18 | 0.14 | 0.16 | 3.02 |
F625W | 0.631 | 12.25-27.0 | 7.53 | 0.22 | 1.23 | 0.20 | 0.05 | 1.27 |
F775W | 0.769 | 11.75-28.5 | 9.54 | 0.89 | 1.35 | 0.22 | 0.14 | 1.63 |
F814W | 0.810 | 11.75-26.5 | 9.13 | 0.65 | 1.08 | 0.18 | 0.05 | 1.27 |
F850LP | 0.903 | 12.25-27.5 | 10.42 | 0.49 | 1.34 | 0.15 | 0.10 | 1.44 |
F105W | 1.05 | 11.75-27.5 | 10.93 | 2.70 | 1.19 | 0.14 | 0.14 | 2.96 |
F110W | 1.15 | 11.75-26.0 | 10.33 | 5.04 | 1.11 | 0.20 | 0.49 | 5.16 |
F125W | 1.25 | 11.75-26.0 | 10.28 | 1.61 | 1.12 | 0.20 | 0.00 | 1.97 |
F140W | 1.39 | 16.0-27.0 | 8.98 | N/A | 0.66 | 1.39 | 0.24 | 1.55 |
F160W | 1.54 | 11.25-26.0 | 10.78 | 1.08 | 1.10 | 0.14 | 0.05 | 1.54 |
Filter | Pivot Wavelength | IGL Value |
F336W | 0.335 | |
F390W | 0.392 | |
F435W | 0.432 | |
F475W | 0.474 | |
F555W | 0.536 | |
F606W | 0.591 | |
F625W | 0.631 | |
F775W | 0.769 | |
F814W | 0.810 | |
F850LP | 0.903 | |
F105W | 1.05 | |
F110W | 1.15 | |
F125W | 1.25 | |
F140W | 1.39 | |
F160W | 1.54 |
6 Comparison With Previous COB Studies
6.1 Comparison to Previous IGL Studies
Here we compare our IGL measurements to previous studies. To make a fair comparison, when comparing to previous IGL works we magnitude zero-point correct our HST number counts to the ground-based ugrizYJH filter system. Over the last decade studies of the IGL have consistently improved due in large part to the accumulation of galaxy number count data reducing cosmic variance uncertainties. Beginning with Driver et al. (2016b), the compilation of number counts from different existing surveys produced IGL measurements with uncertainties between . Cosmic variance contributed a dominant 10% while zero-point uncertainties remained steady at 4.5%. These errors arise from the smaller area coverage of the Galaxy and Mass Assembly Survey (GAMA) (Driver & Robotham, 2010; Liske et al., 2015) relative to WAVES and a significant number of data sets brought to a common magnitude zero-point. In Koushan et al. (2021), the introduction of DEVILS (Davies et al., 2021) increased area coverage at faint magnitudes and thus reduced the number of supplementary data sets required, which reduced cosmic variance and zero-point uncertainties to bring the IGL uncertainty down to 5%. Our extrapolated IGL uncertainties are compared to previous IGL studies in Figure 18 (bottom).
We also compare our IGL measurements to Carter et al. 2025 (submitted), whose work serves as a sister paper to this work and the other capstone paper of the SKYSURF project. Their work introduces a novel technique for measuring the IGL. They measure total galaxy flux in individual fields for galaxies between AB 18-25 mag and extrapolate beyond that range using the ratio between the IGL contained within these bounds and extrapolation to from this work. Given both works originate from the same HST data set, we expect our measurements to be consistent.
Within the respective errors, we find our measurements to be consistent with those of Carter et al. 2025 (submitted), though our IGL values are found to be slightly higher. This can likely be attributed to ProFound’s well documented tendency to extract more flux from faint sources compared to SourceExtractor Robotham et al. (2018).
Band/HST Filter | Pivot Wavelength | IGL Value | Zero-Point | CV | Poisson | Fitting Error | Total Error |
u / F336W | 0.358 | 4.04 | 3.90 | 2.85 | 0.94 | 1.16 | 4.91 |
g / F475W | 0.474 | 5.78 | 0.19 | 0.93 | 0.49 | 2.64 | 2.85 |
r / F606W | 0.631 | 9.14 | 1.79 | 2.51 | 0.90 | 0.42 | 2.84 |
i / F814W | 0.758 | 10.74 | 0.38 | 1.07 | 1.74 | 2.49 | 3.17 |
z / F850LP | 0.890 | 11.24 | 0.23 | 3.06 | 0.37 | 0.18 | 3.09 |
Y / F105W | 1.02 | 11.68 | 1.23 | 1.63 | 0.35 | 0.48 | 2.12 |
J / F125W | 1.25 | 11.13 | 0.82 | 1.12 | 0.36 | 0.50 | 1.52 |
H / F160W | 1.65 | 11.12 | 0.59 | 1.12 | 0.23 | 0.18 | 1.30 |
Band/HST Filter | ABmag Range | Pivot Wavelength | IGL Value | Zero-Point | CV | Poisson | Fitting Error | Total Error |
ABmag | % | % | % | % | % | |||
u / F336W | 14.75-27.0 | 0.358 | 3.30 | 1.72 | 1.25 | 0.18 | 1.16 | 2.42 |
g / F475W | 13.25-26.5 | 0.474 | 4.83 | 0.17 | 1.04 | 0.19 | 2.64 | 2.85 |
r / F606W | 12.25-28.5 | 0.631 | 8.10 | 1.77 | 1.15 | 0.15 | 0.42 | 2.16 |
i / F814W | 11.75-26.0 | 0.758 | 8.91 | 0.35 | 1.11 | 0.15 | 2/39 | 2.66 |
z / F850LP | 12.25-27.5 | 0.890 | 10.39 | 0.25 | 1.30 | 0.15 | 0.18 | 1.34 |
Y / F105W | 11.75-27.5 | 1.02 | 11.20 | 1.18 | 1.20 | 0.18 | 0.48 | 1.76 |
J / F125W | 11.75-27.5 | 1.25 | 10.47 | 0.77 | 1.14 | 0.15 | 0.50 | 1.48 |
H / F160W | 11.25-26.5 | 1.65 | 10.75 | 0.57 | 1.12 | 0.14 | 0.18 | 1.27 |
Band/HST Filter | Pivot Wavelength | IGL Value |
u / F336W | 0.357 | |
g / F475W | 0.474 | |
r / F606W | 0.631 | |
i / F814W | 0.758 | |
z / F850LP | 0.890 | |
Y / F105W | 1.02 | |
J / F125W | 1.25 | |
H / F160W | 1.65 |
The 1.8% error imposed by CV in WAVES in commination with comparable magnitude zero-point uncertainties introduced by using a suite of HST filters limits our improvements over Koushan et al. (2021). Our most precise measurements come from HST filters where we cover significant area and magnitude zero-point corrections are small such as the F814W, F125W, and F160W filters.
From Figure 19 we see that extrapolated IGL measurements are consistently above that of previous works while still within each other’s 1 uncertainties. Our higher IGL values could most likely be attributed to ProFound’s tendency to extract slightly more flux from faint objects than SourceExtractor (Bertin & Arnouts, 1996) as found in Robotham et al. (2018). The HST data used in Driver et al. (2016b); Koushan et al. (2021) was processed using SourceExtractor. In this work, all photometry for DEVILS, WAVES, and HST is provided by ProFound. Nevertheless all measurements are consistent within their quoted errors.
6.2 Comparison to VHE and Direct Measurements of the COB
Finally, we compare our IGL measurements to different measurements of the COB in Figure 19 (right). Our measurements are consistent with estimates of the COB provided by VHE studies (Desai et al., 2017; Fermi-LAT Collaboration et al., 2018; Gréaux et al., 2024). VHE studies provide a probe of the COB independent of the systematic uncertainties discussed in Section 5. Our extrapolated IGL value at is also consistent with the latest direct measurements from New Horizons presented in Postman et al. (2024). Values of , , and from our IGL measurements, the direct measurements of Postman et al. (2024), and VHE measurements of Gréaux et al. (2024) no longer present statistically significant disagreement. Therefore, refining measurements of the COB will provide a better understanding of the remaining foreground flux that constitutes the difference between measurements of the IGL and direct measurements. As the COB and IGL become equivalent, there is merit in further reducing uncertainties in IGL the to in order to serve as constraint on models.
With diffuse and/or unresolved sources of the COB ruled out, the IGL contains all of the information necessary to reconstruct the history of UV-NIR photon production since recombination. Thus the COB can be used as a powerful constraint on models of galaxy formation and evolution as measurements become more precise. Previous works such as (Koushan et al., 2021) demonstrated that the IGL could be used as a constraint on the CSFH. At the time, measurement precision meant that only the normalization of existing CSFH models could be tested. With more precise measurements the shape of the CSFH and models of more complex processes within galaxies can be evaluated using the COB as a constraint.
Finally, we display our measurement of the integrated COB () alongside many different historical measurements, Figure 20. To obtain our measurement, we fit a spline with 7 degrees of freedom to the IGL data from this work, Driver et al. (2016b), and Koushan et al. (2021) between . We then perform the same Monte-Carlo analysis as was done for the IGL measurements discussed in Section 5.1 and integrate 10,000 times. The median, 17th, and 83rd percentiles serve as our measurement of the COB and its 1 uncertainties. We obtain a COB value of , an uncertainty of 2%.
6.3 Prospects For Future Studies of the EBL
We display our COB measurements in context with the current status of the EBL across all wavelengths in Figure 21. Measurements of the COB have finally converged and thus studies of the COB have the potential to become used for precise galactic evolution studies. Current and upcoming facilities such as Euclid, Roman, and SphereX will have the capability to survey much larger areas of sky compared to existing data sets. Euclid has already surveyed a larger area than will ever be accessible to HST. As discussed in Section 5, uncertainties in measurements of the IGL are due to statistical errors caused by limited survey volumes and the need to combine galaxy number counts from multiple facilities. Euclid and the Vera C. Rubin Observatory in particular have the ability to address both of these issues by surveying vast areas at comparable depths to all but the deepest HST fields.
Future surveys will avoid the existing requirement to combine galaxy number counts from different facilities, eliminating magnitude zero-point corrections. The area coverage will also reduce both Poisson and cosmic variance uncertainties to unprecedented levels. However, the HST filter set in combination with existing ugrizYJH data provides a finer sampling of the COB which peaks at .
Studies of the EBL also have potential to be improved outside of the range covered in this work. The successor to SKYSURF, SKYSURFIR (Windhorst et al. (in prep), plans to use archival JWST NIRcam to perform a similar analysis to what has been done in this work. In combination with SphereX, the same level of precision should be accessible for the NIR EBL at .
Outside of the optical and NIR regimes, next generation facilities such as the square kilometer array (SKA) have the potential to unlock new wavelength regimes to precisely study the EBL. The CRB still suffers from the disagreement between upper limits established by direct detection techniques (Fixsen et al., 2011) and the integration of source counts (Gervasi et al., 2008; Vernstrom et al., 2011; Tompkins et al., 2023). The SKA has the potential to rule out significant populations of undetected faint or clustered sources, which had been proposed as a cause of previous disagreements in studies of the COB, as a cause for the current status of radio EBL measurements.
7 Summary
In this work we present the conclusions of the SKYSURF HST legacy archival program in which we produce the most precise measurements of the IGL to date. We produced over 16,000 mosaics starting with calibration level 3 image FLC and FLT image products. We have processed a significant amount of HST data using the same processing pipeline to produce consistent mosaics and object catalogs. Across 16.8 of area we reach completeness limits of AB 24-28.5 mag in different exposure time ranges. By rigorously filtering for problematic fields in biased HST data, we are able to produce reliable galaxy number counts across 17 of 22 HST filters from which mosaics were constructed. We use our HST number counts alongside WAVES and DEVILS to measure the IGL at 17 different wavelengths between 0.3-1.6 m by magnitude zero-point correcting the WAVES and DEVILS number counts to the HST filter set. We then extensively discuss the significant sources of error affecting the IGL, Poisson, cosmic variance, and magnitude zero point uncertainties. We then comment on the self-consistency of our HST IGL measurements in adjacent HST filters.
We also select 8 HST filters corrected to the ugrizYJH filter set used in WAVES to compare with existing literature which has done the same. In both cases, we report extrapolated IGL values and lower limits obtained by integrating across the magnitude range covered by the data. We then place our 8 IGL measurements on the ugrizYJH filter set in context with previous IGL studies and measurements of the COB at large. We also quote measurements at the mean value between our lower limits and extrapolated IGL which serve as the most conservative result in this work.
8 Acknowledgments
This work was made possible by the following grants, HST programs AR-09955 and AR-15810 provided by NASA through the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Timothy Carleton is thankful for support from the Beus Center for Cosmic Foundations.
9 Software